OpenStudy (anonymous):

Solve the equation for W: ln abs(W-300)=(1/25)t+ln(1100).

OpenStudy (anonymous):

\[\ln|W-300|=\frac{1}{25}t+\ln1100\] To rewrite, I'll use the fact that \(\ln e=1\): \[\ln|W-300|=\frac{1}{25}t\ln e+\ln1100\] Another property, \(b\ln a=\ln (a^b)\): \[\ln|W-300|=\ln e^{1/25~t}+\ln1100\] Another property, \(\ln a+\ln b=\ln (ab)\): \[\ln|W-300|=\ln 1100e^{1/25~t}\] Now you can take the exponential of both sides: \[e^{\ln|W-300|}=e^{\ln 1100e^{1/25~t}}\] which reduces to the simpler equation, \[|W-300|=1100e^{1/25~t}\]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

yw